OpenStudy (anonymous):

The bird population on an island is declining at a rate of 2.2% per year. If the population was 3500 in the year 2009, which is the best prediction of the population in the year 2014? (Points : 3) 2730 3062 3132 3423 I think the answer is 3423 am i right?

OpenStudy (anonymous):

@iambatman @danish071996

OpenStudy (freckles):

What formula are you using?

OpenStudy (anonymous):

i figured it out could you help with a differen one? @freckles

OpenStudy (anonymous):

Which expression is equivalent to ? (Points : 3) Option A: 2 Option B: 16 Option C: ^3 squared 4 Option D: ^6 squared 16

OpenStudy (freckles):

Your guess is good as mine on that one. Sounds like you just have to guess since they didn't given you an expression to use.

OpenStudy (anonymous):

oops they did i forgot to put it its 4^1/3 times 4 ^1/6

OpenStudy (freckles):

add the exponents since you have multiplication can you do 1/3+1/6?

OpenStudy (anonymous):

3/6

OpenStudy (freckles):

\[\frac{1}{3}+\frac{1}{6} \\ \text{ ..} \frac{1}{3}=\frac{1 \cdot 2 }{ 3 \cdot 2 }=\frac{2}{6} \\ \frac{2}{6}+\frac{1}{6}=\frac{3}{6}=\frac{1}{2} \] great so you have 4^(1/2)

OpenStudy (freckles):

do you know the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (freckles):

\[4^{\frac{1}{3}+\frac{1}{6}}=4^\frac{3}{6}=4^\frac{1}{2}=\sqrt{4}=2 \] :) good

OpenStudy (anonymous):

thats our answer?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

I don't think we can simplify 2 any further . 2 is just 2 .